What is the ratio of De Brogile's wave lengths of Mg2+ ion and C+ ion accelerated with 100 volts potentials each.


1
Expert's answer
2021-07-28T01:22:07-0400

"\lambda=\frac{h}{p}=\frac{h}{\sqrt{2mE}}
\frac{h}{\sqrt{2mqV}}"


"\lambda=\frac{6.6×10^{-34}}{\sqrt{2×9.1×10^{-31}×1.6×10^{-19}×100}}"


"\lambda=8.6×10^{-11}"